Amanda Fyles and Maria Palacios – The Damage of Water Run-Off

Your presentation to the class was excellent and included a substantial amount of material that went beyond what the other “run-off groups” did. The techniques suggested by the project at Virginia Tech were excellent include. They show that there are practical approaches to reducing runoff issues, and that other college and university campuses are already making use of them(!) The paper is also excellent for the most part, although there were a few small writing and technical issues here and there. Also see the specific comments below.

Specific Comments

1) “The liver of our existence?” I think I know what you meant, but this is somewhat strange-sounding as the first sentence(!) I think you could have said something like: “Soils perform some of the same functions for terrestrial ecosystems that the liver performs in our bodies.”

2) The technical name for that distance measuring device is the odometer. It works by the same principle as the mechnical odometers found in older automobiles. As the wheel travels over the ground, a train of gears turns the gauge showing the distance traveled.

3) You should say somewhere that you were responsible for Region 3. I think your group did by far the best and most accurate job of actually carrying out the measurements, and your initiative in finding the large-scale paper campus map at Buildings and Grounds was a big plus for your project!

4) To go ahead with this sort of project and make a more complete/accurate survey of the campus, I actually think now that normalizing to get a percentage of each type of the surface within each small region (“block”) would be the best way to proceed only if all of the blocks had roughly the same total area. There might be a good reason to use the actual areas (not percents) since the runoff water amount is going to be roughly proportional to the actual area for each type of surface.

5) “Suggests a evel of legitimacy” sounds harsh here – maybe say “which suggests that the categorization and weighting of the data was meaningful.”

6) Careful, there is only one SD for the whole data set. You use the formula z = \frac{x−\overline{x}}{SD} on each measurement to get a z-score.

7) This came out wrong. I am sure you meant to say that the run-off problem is much more severe today than that seen fifty or one hundred years ago.

8) I’m not sure that you want to claim this approach “maximizes land use efficiency.” You have not established a criterion for efficiency and it would be easy to argue that a more “efficient” approach would be to pack as many parking spaces in as possible (if efficient means maximizing parking capacity within a given area). So it would be better to say something like “This would address the need for parking and decrease the contribution of parking areas to run-off.”

Final Project Grades Presentation: A (95) Paper Content: A (95) Paper Form: A- (92)